### v1

# Assessing the impact of variance heterogeneity and misspecification in mixed-effects location-scale models

Abstract
Purpose: Linear Mixed Model (LMM) is a common statistical approach to model the relation between exposure and outcome while capturing individual variability through random effects. However, this model assumes the homogeneity of the error term’s variance. Breaking this assumption, known as homoscedasticity, can bias estimates and, consequently, may change a study’s conclusions. If this assumption is unmet, the mixed-effect location-scale model (MELSM) offers a solution to account for within-individual variability.
Methods: Our work explores how LMMs and MELSMs behave when the homoscedasticity assumption is not met. Further, we study how misspecification affects inference for MELSM. To this aim, we propose a simulation study with longitudinal data and evaluate the estimates’ bias and coverage.
Results: Our simulations show that neglecting heteroscedasticity in LMMs leads to loss of coverage for the estimated coefficients and biases the estimates of the standard deviations of the random effects. In MELSMs, scale misspecification does not bias the location model, but location misspecification alters the scale estimates.
Conclusion: Our simulation study illustrates the importance of modelling heteroscedasticity, with potential implications beyond mixed effect models, for generalised linear mixed models for non-normal outcomes and joint models with survival data.
keywords:
Mixed-effects location-scale models (MELSM), Heteroscedasticity, Model misspecification, Distributional regression, Simulation study
1 Background
Longitudinal studies are critical to answering research questions in many fields because of their ability to highlight trends and changes over time. In medicine, these analyses are fundamental to capturing a disease’s evolution through different stages of life, as well as measuring the changes in treatments’ effects. In social sciences, they are used to track the evolution of socioeconomic phenomena, such as the impact of where one lives on economic outcomes [22], and the associated impact of policies on these outcomes.
In the analysis of longitudinal data, linear mixed models (LMMs) emerged as a popular option to model repeated measures for the same statistical unit, such as individuals or countries, because of their interpretability and the availability of implementations in multiple statistical software [18, 3]. The key advantage of LMMs is that one can model the population average as a function of time and covariates – these population-level relations are known as fixed effects –, and individual-specific predictions, known as random effects, which capture the correlation between the measures for the same unit and quantify the deviation with respect to the population average. Random effects are assumed to be drawn from a common distribution (usually normal) with zero mean.
An underlying assumption of LMMs is homoscedasticity: the residual variance is assumed constant across all observations and covariates. In other words, the effect of unobserved covariates, which is captured by the error term, contributes to variation in equal ways for all units. While some options to validate this assumption exist (e.g.L̃evene’s test [8] for homogeneity of variance, or fitted-residual plots to graphically assess the absence of unexplained patterns), these are not commonly reported in practice. While heteroscedasticity does not necessarily bias fixed effect estimates, ignoring this heterogeneity can impact variance estimates and, consequently, result in Type I and II errors [13, 14].
Modelling heterogeneity may even reflect important population and individual characteristics. For instance, in cardiovascular research, the notion of blood pressure variability [15] is becoming more commonly used as a risk factor for stroke and other events, as large fluctuations over time might be indicative of arterial health and other issues. This could lead to finding new treatments which do not only bring blood pressure to a normal range but also stabilise it around some average values.
In the statistics literature, multiple models model the variance as well as the mean. Hedeker et al [4, 5] introduced a mixed-effect location-scale model (MELSM) to model both the location parameter (the mean) and the scale parameter (the residual standard deviation) as a function of covariates and random effects. The model was initially applied to quantify the within-subject heterogeneity in ecological momentary assessments to study mood variation [11], but then has been applied to other domains, e.g. cystic fibrosis studies [9]. A general modelling framework is provided in gamlss [12] and its Bayesian version (bamlss, 17) with a large range of distributions, all parameters can be modelled using fixed and random effects (although not many examples explicitly model within-individual variability using random effects). All these options fall under the umbrella of distributional regression models, where the conditional distribution of the outcome given the covariates is studied via its parameters.
Despite the existence of such solutions, modelling heteroscedasticity is not often done in practice. Walters et al [20] discussed how software availability and a lack of best practices in specifying the scale model may explain this lack of adoption. With recent user-friendly software developments such as [2] and [12], only the second barrier remains. Further, methodological challenges exist when using MELSMs. For example, although the problem of power analysis and testing the homogeneity of variance in MELSM was the object of some recent studies [20, 21], the relationship between location and scale is not fully understood, and how the misspecification of one affects the other.
Our work aims to fill this gap by exploring how common modelling practices and associated misspecifications impact the estimates of interest. Through a comprehensive set of simulations, we explore how ignoring heteroscedasticity and misspecifying the location and scale models impact random and fixed effects estimations. Our work highlights how MELSM offers a robust framework for handling heteroscedasticity.
The paper is structured as follows. Section 2 introduces our notations and the studied MELSM. Section 3 describes the simulation setting. In Section 4, we study six common practices in specifying the scale model and explore how misspecifications affect MELSM’s estimates. Finally, in Section 5, we discuss the insights from the simulations, their limitations, and potential further developments.
2 Mixed-effect location-scale model
Consider the normally-distributed outcome $y_{i,j}\in\mathbb{R}$ for subject $i\in[1,N]$ at encounter $j\in[1,n_{i}]$. A linear mixed effect model assumes the following relation between outcome and covariates:
$y_{i,j}=x_{i,j}^{yT}\beta^{y}+z_{i,j}^{yT}u^{y}_{i}+\epsilon_{i,j}$
A linear mixed-effect model assumes that $y_{i,j}\sim\mathcal{N}(x_{i,j}^{yT}\beta^{y}+z_{i,j}^{yT}u^{y}_{i},\omega^{2})$, where the standard deviation $\omega$ is constant for all observations (homoscedasticity). The MELSM relaxes this assumption by specifying a model for the scale, which can depend on random effects, as well as covariates, i.e.:
$\omega_{i,j}:=\exp(x_{i,j}^{\omega T}\beta^{\omega}+z_{i,j}^{\omega T}u^{\omega}_{i})$
where $\beta^{\omega}$ gives some insights on the contribution of the different covariates on the variance. We model the logarithm of the standard deviation to ensure non-negativity. The random effect vectors $u^{y}_{i}$ and $u^{\omega}_{i}$ are to be interpreted as subject-specific deviations from the population mean and standard deviation, respectively. For ease of notation, random intercepts $u^{y}_{0,i}$ and $u^{\omega}_{0,i}$ are implicitly included by setting the first column of the corresponding covariate matrices to a vector of 1s, while random slopes $u^{y}_{1,i}$ and $u^{\omega}_{1,i}$ are obtained by setting the column to the vector of observation times. The random effects for the location and scale models are assumed to be normally distributed with mean zero, and a correlation between random effects can be estimated. For example, in the simple setting with just a random intercept for both the location and scale models, we can specify the correlation $\rho$ to capture the relationship between the two models:
$\begin{pmatrix}u_{i}^{y}\\ u_{i}^{w}\end{pmatrix}\sim\mathcal{N}\begin{bmatrix}\begin{pmatrix}0\\ 0\end{pmatrix}\!\!,\begin{pmatrix}\sigma^{y^{2}}&\rho\sigma^{y}\sigma^{\omega}\\ \rho\sigma^{y}\sigma^{\omega}&\sigma^{\omega^{2}}\end{pmatrix}\end{bmatrix}$ (1)
We denote the covariance matrix in (1) as $\Sigma$.
3 Simulation study
This section describes the data generation and evaluation used in the proposed simulations to evaluate the robustness of MELSM to different misspecifications.
3.1 Data generation
We generate simulation data based on the Primary Biliary Cholangitis (PBC) dataset, publicly available in the R package survival, commonly used as a test for longitudinal modelling. The dataset results from a clinical trial at Mayo Clinic studying the impact of D-penicillamine on PBC, an auto-immune condition affecting the liver. From the dataset, we extract four covariates at baseline: age, albumin, triglycerides (trig), and platelet.
Using these covariates, we generate a synthetic population of $N$ individuals by drawing from a normal distribution with the covariance defined by the previously extracted and standardised covariates at baseline. Note that $N$ is a parameter of the simulation and not the observed PBC population size, as one of our aims is to study how an increasing number of patients impacts our conclusions. Formally, we draw $x_{i,0}$ – where $i$ describes the index associate with the patient and 00, the baseline encounter – from a multivariate normal distribution $\text{MVN}(0,\Sigma)$ with $\Sigma$, the covariance matrix described in Table 1.
age albumin trig platelet
age 1.02 -0.23 0.02 -0.14
albumin -0.23 0.90 -0.10 0.18
trig 0.02 -0.10 1.00 0.10
platelet -0.14 0.18 0.10 0.89
Table 1: Observed covariance matrix for baseline covariates in the PBC dataset.
As we model longitudinal data, we then draw a number of encounters $J_{i}$ from a discrete uniform distribution for each patient $i$:
$J_{i}\sim\mathcal{U}\{1,2M-1\}$
where $M$ is a parameter of the simulation study representing the average number of encounters per subject across the population. For each encounter, we compute the age at this encounter as follows:
$\text{age}_{i,j}=\text{age}_{i,0}+j$
and we use this covariate as the only time-varying variable while all other covariates remain constant over time. Unless otherwise specified, we consider a total of $N=200$ patients with an average number of observations $M=15$. Further, we simulate random intercepts for all individuals defined as in Equation 1. We set $\rho=0$, $\sigma^{y}=2$ and $\sigma^{\omega}=1$.
MELSM models a longitudinal outcome of interest. To this end, we generate the outcome as follows:
$y_{i,j}$ $=\beta^{y}\cdot x_{i,j}+u^{y}_{0,i}+\epsilon_{i,j}$ (Location)
$\epsilon_{i,j}$ $\sim\mathcal{N}(0,\omega^{2}_{i,j})$
$\omega_{i,j}$ $=\exp(\beta^{\omega}\cdot x_{i,j}+u^{\omega}_{0,i})$ (Scale)
In our simulations, we set the fixed effect to depend on two covariates for location and scale (to allow for a variety of plausible misspecification types), as summarised in Table 2.
age albumin trig platelet
$\beta^{y}$ 0.5 0.5 0 0
$\beta^{\omega}$ 0.8 0 0.8 0
Table 2: Fixed effects for location and scale.
3.2 Estimands of interest
Our simulations explore the recovery of the fixed and random effects for both scale and location models. To this end, we report the box plots and the associated coverage (for 95% credible intervals) for the estimated coefficients of interest across 100 repetitions of the simulation described in Section 3.1. Formally, the coverage of the parameter $\theta$ (in percent) is defined as follows:
$\sum_{s}\mathbbm{1}(L(\hat{\theta},\alpha)<\theta<U(\hat{\theta},\alpha))$
in which $L$ and $U$ are the lower and upper bounds of the estimated $\hat{\theta}$ at the confidence level $\alpha$.
3.3 Software
All experiments are implemented in R with code made available on GitHub [^1] https://github.com/Jeanselme/IndividualVariability. We fit the MELSM models using the brms package [2], with default priors, two chains, and a burn-in of 1000 iterations for a total of 2000 iterations.
4 Results
This section describes common practices when using mixed effect models and their impact on parameters’ estimates. For each model, we report the formula in pseudo-R code for reproducibility.
4.1 Practice 1: Ignoring heteroscedasticity
By relying on LMM, practitioners typically make an assumption of homoscedasticity, ignoring potential subject-specific differences in variance. In this first setting, we compare an LMM, assuming a constant scale, with the properly specified MELSM model. The two models share the same (correct) specification for the location, with only the scale differing:
  • MELSM correctly models the scale.
            y ~ age + albumin + (1|id)
            log(omega) ~ age + trig + (1|id)
  • LMM assumes homoscedasticity.
            y ~  age + albumin + (1|id)
            log(omega) ~ 1
In this first set of simulations, we explore the impact of varying the number of patients and observations per patients, i.e. $N\in\{100,300,500,1000\}$ and $M\in\{5,10,20\}$. These values were chosen to approximate different data amounts and frequencies as common in real longitudinal data settings. Figure 1 presents the estimated fixed effects associated with age in the location and scale models under varying numbers of individuals $N$ across 100 simulations. Similarly, Figure 2 presents the same results under varying average numbers of observations per patients. In each figure, the red vertical line indicates the true estimand’s value.
Figure 1: Estimated age effect on location (left panel) and scale (right panel) for increasing number of individuals. For each model, coverage is reported as a percentage next to the corresponding boxplot.
Figure 2: Estimated age effect on location (left panel) and scale (right panel) for increasing average number of observations per individual. For each model, coverage is reported as a percentage next to the corresponding boxplot.
Scale misspecification impacts the precision of location estimates.
Under both analyses, Figures 1 and 2 show that the estimated age coefficients in the location model, although centered around the true value, display larger variance for LMM than MELSM across all sizes of the generated datasets.
Coverage is low in misspecified settings.
The coverage in the LMM is well below the nominal level of 95%, indicating that the standard errors are underestimated. Consequently, the misspecification of the scale has a direct impact on the location.
Larger number of individuals and encounters reduce error.
Both models benefit from a larger number of individuals and encounters as shown by reduced variance in the location estimates, and reduced variance for MELSM scale estimates.
4.2 Practice 2: Misspecifying location and scale in MELSM
As the parametric form of either scale or location is rarely known in advance, we explore the influence of different types of misspecification (adding irrelevant covariates, neglecting relevant covariates or misspecifying the random effect structure) within MELSM. Specifically, we consider the following modelling of the generated data:
  • Correct is the correctly specified model.
            y ~ age + albumin + (1|id),
            log(omega) ~ age + trig + (1|id)
  • All considers all covariates for both models.
            y ~ age + albumin + trig + platelet + (1|id)
            log(omega) ~ age + albumin + trig + platelet + (1|id)
  • Misspecified $y$ ignores age in the location model.
            y ~ albumin + (1|id)
            log(omega) ~ age + trig + (1|id)
  • Misspecified $\omega$ does not consider age in the scale expression.
            y ~ age + albumin + (1|id)
            log(omega) ~ trig + (1|id)
  • No $u^{\omega}$ ignores the random effect associated with the scale model.
            y ~ age + albumin + (1|id)
            log(omega) ~ age + trig
  • No $u^{\omega}$ and misspecified $\omega$ combines the two previous scenarios, with age and random intercept ignored in the scale model.
            y ~ age + albumin + (1|id)
            log(omega) ~ trig
Figure 3: Estimated age effect for location (left panel) and scale (right panel). For each model, coverage is reported as percentage next to the corresponding boxplot.
Figure 4: Estimated standard deviation of random intercepts for location (left panel) and scale (right panel). For each model, coverage is reported as percentage next to the corresponding boxplot.
Figure 3 presents the fixed effect associated with the age covariate, while Figure 4 describes the estimated standard deviation of the random intercept across the simulations. On the right, we display the coverage in percent over the different simulations. For consistency, we display all method specifications in these figures. Note that some do not estimate all quantities, hence the absence of some estimates.
Adding superfluous parameters does not invalidate inference.
In all figures, the correctly specified model (Correct) and the one considering all covariates (All) result in similar estimates. This shows that, with enough data, all models that include the correct covariates as a special case will lead to a correct inference. Note that, however, considering all covariates results in an increased number of parameters and associated computational time — the method considering all covariates is more than 50% slower than the correctly specified model.
Not all scale misspecification types affect the location in the same way.
The age effect in the location model is well estimated across all models, although the variance of the estimates increases as the scale model departs from the true one (as observed with LMM). The largest variance is observed for No $u^{\omega}$ + Mis. $\omega$, where both a relevant covariate and the random intercept are missing in the scale model. The same model also shows, on average, lower standard deviations for the scale random intercept than the true value.
Location misspecification impacts scale.
In the misspecified location model Mis. $y$, the scale fixed effect estimate and random intercept are biased, despite presenting the right scale specification. As the location parameterisation does not capture the full variation of the data, the residual variance is different from the true value, and therefore, all scale estimates are biased (Appendix A provides a theoretical intuition of why this occurs).
However, the opposite only affects the variance of the estimates: when the scale is misspecified, as in Mis. $\omega$, the location’s age effect estimate remains unbiased. When the scale model ignores age (Mis. $\omega$) or ignores the random intercept (No $u^{\omega}$), the location effect presents a larger variance. Further, when both random intercept and misspecification of the fixed effect occur (No $u^{\omega}$ + Mis. $\omega$), the variance more than doubles.
Therefore, location misspecification results in biased scale estimates, while the opposite only impacts the variance of the estimands.
There is a relationship between random effects and residual variance.
The estimated standard deviation of location random intercept can incur bias when the scale model is incorrectly specified (and therefore, the same happens in LMM when there is heteroscedasticity between the individuals). The standard deviation for the random effect of the scale model is biased in all considered scenarios unless the fixed effect structure for both location and scale are correctly specified.
4.3 Practice 3: Ignoring non-linear time trend in location model
Another common type of misspecification in everyday practice is represented by the choice of a linear relationship between the outcome mean and the time variable when actually a non-linear trend is present. We, therefore, explore how non-linearity would affect the estimates of interest. To this end, we alter the data generation mechanism by applying a sinusoidal transformation on age as follows:
$y_{i}=\beta^{y}_{age}\sin(age_{i})+\beta^{y}_{alb}albumin_{i}+u^{y}_{i}+\epsilon_{i}$
while we keep a linear relationship with the scale parameter. In this context, we compare the properly specified models with the common misspecification ignoring the non-linearity (while the scale is correctly specified):
  • Correct consists of the properly specified model under this data generation process.
              y ~ sin(age) + albumin + (1|id),
              log(omega) ~ age + trig + (1|id)
  • Non sinus ignores the non-linearity on the covariates in the location.
              y ~  age + albumin + (1|id)
              log(omega) ~ age + trig + (1|id)
Figure 5 presents the fixed effect estimate associated with the age covariate.
Figure 5: Estimated age effect for location (left panel) and scale (right panel). For each model, coverage is reported as percentage next to the corresponding boxplot.
Non-linearity biases location fixed effects.
In this case, the misspecification of the functional form of age leads to a biased estimate of the effect of the covariate on the location model. As the scale relies on age and not a functional transformation, the scale model remains non-impacted.
4.4 Practice 4: Misspecifying the random effect structure
One drawback of MELSM is that the computation becomes heavier as the number of random effects increases. For this reason, a practitioner might decide to have a more parsimonious model structure with only random intercepts for both location and scale models, even if random slopes might actually be justified in the analysis.
In the simulation settings discussed so far, we have generated the data assuming only a random intercept for the location and a random intercept for the scale, while in this setting, we alter the data generative process by adding an uncorrelated random slope for both location and scale. We then explore how different misspecifications may alter the estimates of the fixed effect. In these experiments, the generative process of the scale and location is as follows:
$y_{i,j}$ $=\beta^{y}\cdot x_{i,j}+u^{y}_{0,i}+\bm{u^{y}_{age,i}\text{age}_{i,j}}+\epsilon_{i,j}$
$\epsilon_{i,j}$ $\sim\mathcal{N}(0,\omega^{2}_{i,j})$
$\omega_{i,j}$ $=\exp(\beta^{\omega}\cdot x_{i,j}+u^{\omega}_{0,i}+\bm{u^{\omega}_{age,i}\text{age}_{i,j}})$
in which both $u^{y}_{age,i}$ and $u^{\omega}_{age,i}$ are drawn from normal distributions: $\mathcal{N}(0,1)$ resp. $\mathcal{N}(0,0.5^{2})$. In this context, we consider three models that properly control on the correct covariates but specify different random effects:
  • Correct models both random slopes.
              y ~ age + albumin + (1 + age|id),
              log(omega) ~ age + trig + (1 + age|id)
  • No $u_{age}^{\omega}$ ignores the scale random slope.
              y ~ age + albumin + (1 + age|id),
              log(omega) ~ age + trig + (1|id)
  • No slopes ignores both random slopes.
              y ~  age + albumin + (1|id),
              log(omega) ~ age + trig + (1|id)
Figure 6 presents the fixed effect estimate associated with the age covariate, and Figure 7, the estimated random intercepts for all models.
Figure 6: Estimated age effect for location (left panel) and scale (right panel). For each model, coverage is reported as percentage next to the corresponding boxplot.
Figure 7: Estimated standard deviation of random intercepts for location (left panel) and scale (right panel). For each model, coverage is reported as percentage next to the corresponding boxplot.
Missing random slope will mostly affect the corresponding time fixed effect.
When a non-zero random slope is neglected, the estimates of the age effect (that corresponds to the fixed slope or population effect) are centred around the true value, but the coverage is much lower than the nominal level. It is worth mentioning that missing the random slope in the scale model does not influence the fixed effect in the location model, but in the opposite case, the age effect in the scale model shows a loss of coverage. When random intercepts and slopes are independent, neglecting the random slopes induces no changes in the standard deviations of the random intercepts.
4.5 Practice 5: Misspecifying random effect distributions
Another common practice is to only consider Gaussian distributions for the random intercepts. In this scenario, we explore the misspecification of the random intercept distribution by sampling the random intercept from Student distributions with 3 degrees of freedom instead of a normal distribution, and compare the two following models:
  • Student correctly specifies the random intercept distributions.
              y ~  age + albumin + (1|gr(id, dist=‘student’))
              log(omega) ~ age + trig + (1|gr(id, dist=‘student’))
  • Gaussian assumes Gaussian distributions (the default option).
              y ~ age + albumin + (1|id),
              log(omega) ~ age + trig + (1|id)
This setting resulted in no difference in the fixed effect estimates, our analysis therefore focuses on the random intercepts in Figure 8.
Figure 8: Estimated standard deviation of random intercepts for location (left panel) and scale (right panel). For each model, coverage is reported as percentage next to the corresponding boxplot.
Figure 9: Cumulative distribution function of random effects for the location (left panel) and scale (right panel). The standard deviations of the distributions correspond to the medians in the previous figure.
Misspecified distributions do not impact on random effect estimation.
When the true random effect distribution is t-Student, the standard deviation of the Gaussian distribution will be estimated so that the Gaussian distribution matches the original distribution, as illustrated in Figure 9. This results in similar estimated distributions despite misestimated standard deviations. This result is consistent with the LMM literature (e.g.16).
5 Discussion
In this work, we have presented a set of simulations to study the impact of misspecification on mixed-effect location-scale models, with a focus on modelling heteroscedasticity as opposed to the assumption of constant residual variance made by standard linear mixed models.
Our simulations offer valuable insights into MELSM. First, scale misspecification increases the variance of the location estimates but does not introduce bias. In contrast, location misspecification biases scale estimates. Therefore, all conclusions about covariate effects in the scale model are to be referred to the specific set of covariates and functional forms used in the location. Further, there is an interplay between the random effects in the location model and the residual variance. Different models for residual variance influence the standard deviations of the location random effects. Finally, misspecification of the model structure, such as neglecting a relevant covariate, leads to larger bias than misspecification of the functional form, as including relevant covariates helps account for part of their effects, even if the functional form is not correct. These insights show the critical importance of modelling residual variance, inviting statisticians to model this quantity even under imperfect knowledge of its structure.
As a simulation, our findings have limitations. While we have used a general set of parameters to improve the generalisability of our conclusions, more extreme parameter settings could depart from the results shown here. Similarly, our simulations rely on two random effects and do not consider random effects for B-spline coefficients. Our conclusion on the misspecification of random effect distribution considers a symmetric alternative to the normal distribution and not more skewed distributions, which could impact this finding. Finally, our work focuses on the distribution shape and not on the prediction of individual random effects, which the linear mixed models literature suggests being more affected by the misspecification (see 7, 6, 19 for reviews on the topic).
This work opens avenues for evaluating misspecifications in other location-scale models — for instance, the addition of time-varying or non-normally distributed covariates. Beyond the presented setting, MELSMs are often used as part of more complex models. Particularly, the study of joint models with within-individual variability [1, 10] would provide insights into the relationship between the misspecification of the longitudinal model and the association with the hazard in time-to-event data analysis.
Declarations
For the purpose of open access, the author has applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising from this submission.
Ethics and Consent to Participate declarations. Not applicable.
Consent for publication. Not applicable.
Availability of data and materials. All data generated and code to reproduce the current study are available at https://github.com/Jeanselme/IndividualVariability.
Competing interests. JKB has received research funding for unrelated work from F. Hoffmann-La Roche Ltd.
Funding. This work was supported by the UK Medical Research Council grant “Looking beyond the mean: what within-person variability can tell us about dementia, cardiovascular disease and cystic fibrosis” (MR/V020595/1). JKB was supported through the United Kingdom Medical Research Council programme (grant MC_UU_00002/5) and theme (grant MC_UU_00040/02 – Precision Medicine) funding. MP’s research is supported by the Ulverscroft Vision Research Group (UCL).
Authors’ contributions. VJ: Formal analysis, software, writing—original draft. JKB: Funding acquisition. MP, JKB: Supervision. VJ, MP, JKB: Conceptualization, writing—reviewing and editing.
Appendix A Why do we expect a biased location model to have a larger scale error?
Let us simplify the problem by considering the following generative process:
$Y=\beta X+\epsilon$
with $\epsilon\sim\mathcal{N}(0,\omega)$.
A biased scale model can be expressed as:
$Y=(\beta+\delta)X+\epsilon^{\prime}$
with $\epsilon^{\prime}\sim\mathcal{N}(0,\omega^{\prime})$ and $\delta$ the bias from the true coefficient.
Computing the difference between the generative process and the biased model results in:
$\epsilon^{\prime}=\epsilon-\delta X$
with $X\sim MVN(0,\Sigma)$. And therefore, one can express the variance relation as:
$\omega^{\prime}=\omega+\delta^{T}\Sigma\delta$
Note that the exact quantification of these quantities are dependent upon the correlation structure between covariates. However, as $\Sigma$ is positive semi-definite:
$\omega^{\prime}\geq\omega$
This means that the estimated variance of a biased model is larger than the variance of the data.